OpenStudy (cheesecakekitten):

Marcus is eating a meal that is made up of an entrée that is 45 calories per ounce and a side dish that is 30 calories per ounce. He wants his meal to total 510 calories. Let e represent the number of ounces of the entrée that he eats, and let s represent the number of ounces of the side dish that he eats. Which function relates the number of ounces of the side dish that he eats to the number of ounces of the entrée that he eats?

OpenStudy (cheesecakekitten):

Will give medal and fan!!!!

OpenStudy (anonymous):

510 = 45e + 30s Solve for s.

OpenStudy (anonymous):

typo 30s = -45e + 510

OpenStudy (cheesecakekitten):

i don't know how to solve for variables

OpenStudy (anonymous):

is it mulitiple choice?

OpenStudy (cheesecakekitten):

yes

OpenStudy (anonymous):

what are the choices?

OpenStudy (cheesecakekitten):

i will give out answers brb

OpenStudy (cheesecakekitten):

A. s = 2/3e - 17 B. s = 3/2e - 11 1/3 C. s = -2/3e + 11 1/3 D. s = -3/2e + 17

OpenStudy (cheesecakekitten):

srry it took so long

OpenStudy (anonymous):

hold on, this question is already asked haha math.openstudy.com/updates/546cfb25e4b0db9322961c63

OpenStudy (cheesecakekitten):

there is no answer

OpenStudy (anonymous):

ok well lets solve this then. the answers seem to solve for S right? if you want to find s, you need to get it by itself. so to get it by itself, you have to devide s by 30 since its being multiplied to cancel it out. so on the right side, devide each term by 30

OpenStudy (cheesecakekitten):

oke i will just a sec

OpenStudy (cheesecakekitten):

oke i have no idea how to do this i'm not good at equations

OpenStudy (anonymous):

30s = -45e + 510 so this is the set up equasion were going to use. To get a variable by itself you have to do the opposite operation the variable contains. see how the number is right away followed by s? that means they are being multiplied. so if you want to cancel the thirty, you have to divide everything on the left side by 30, which would cancel out the 30 on the 30s, leaving it just S. got that?

OpenStudy (cheesecakekitten):

so s is 15.5?

OpenStudy (anonymous):

no S is going to be the end of the equasion when you solve. So to cancel thirty S you divide everything by 30. so 30s divided by 30= 1s, or just s. So the first thing would be S= then you divide -45e by 30. the answer would be -1.5, but in this case we would put it as a fraction. so -3/2. so now we have S=-3/2. for the last thing you have to devide 510 by 30 also. which is seventeen. so the anwer would be S=-3/2+17

OpenStudy (cheesecakekitten):

oh thanks

OpenStudy (anonymous):

yup :)

OpenStudy (cheesecakekitten):

medal and fan as said

OpenStudy (cheesecakekitten):

I'm your first fan! XD

OpenStudy (anonymous):

yup haha im just getting started

OpenStudy (cheesecakekitten):

i'm fairly new here as well